# Reeb spaces of smooth functions associated to globally similar graphs of smooth functions

### 12 Mar 2026 v2

@@ Abstract.
Here, the Reeb space of a smooth function on a manifold with no boundary is fundamental and important. This is the naturally topologized quotient space of the manifold, consisting of all connected components (contours) of the function and is a graph under a certain nice situation. Related studies related to the present study were started due to interest of the author in theory of Reeb spaces of so-called non-proper functions. For proper functions, in 2020s, related studies have developed mainly due to Gelbukh and Saeki.
@@
  • $K(C_{p})-C_{p}$ does not contain any point of the form $(c_{i}(x),x,{(0)}_{j=1}^{m-1})$ ($x\in S(c_{i})$).
@@
A smooth function $c_{p,0}:\mathbb{R}\rightarrow\mathbb{R}$ defined by $c_{p,0}(x):=\frac{2+\sin(e^{x})}{p(x)}$ enjoys the following properties.
@@
    ${c_{p,0}}^{\prime}(x)=\frac{e^{x}\times\cos(e^{x})\times{p(x)}-(2+\sin(e^{x}))\times p^{\prime}(x)}{{p(x)}^{2}}$.
This is a kind of fundamental exercises from elementary calculus. By considering $x$ satisfying $\cos(e^{x^{2}})=\pm 1$ and asymptotic behaviors of $e^{x}$ and real polynomial functions (or so-called real algebraic functions), we have the following.
@@
Let $c_{p,0}$ be a function of Proposition 1. There exist a sequence $\{r_{{\rm P},j}\}_{j=1}^{\infty}$ of positive numbers diverging to $\infty$ such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm P},j})\}_{j=1}^{\infty}$ diverges to $\infty$ and one such that the sequence $\{{c_{0}}^{\prime}(r_{{\rm P},j})\}_{j=1}^{\infty}$ diverges to -$\infty$. There exist a sequence $\{r_{{\rm N},j}\}_{j=1}^{\infty}$ of negative numbers diverging to -$\infty$ such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm N},j})\}_{j=1}^{\infty}$ diverges to $\infty$ and one such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm N},j})\}_{j=1}^{\infty}$ diverges to -$\infty$.
@@
We can define $c_{1}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})$ and $c_{2}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+c_{p,0}(x_{1,2})$, with $c_{p,0}$ from Proposition 1, and have a case for the former statement. More precisely, we choose $p(x)=x^{2}+1$ for example.
We can define $c_{1}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+\frac{1}{2}\times c_{p,0}(x_{1,2})$ and $c_{2}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+c_{p,0}(x_{1,2})$, with $c_{p,0}$ from Proposition 1, and have a case for the latter statement.
@@
We remember Propositions 1 and 2. we define a function of a similar type we can use in the present situation. In short, we define $c_{e}(x):=\frac{2+\sin(e^{x^{4}})}{e^{x^{2}}}$ instead. The values of this function are always positive and converge to 0 at each infinity. Its 1st derivative is calculated as ${c_{e}}^{\prime}(x)=\frac{(4x^{3}e^{x^{4}}\cos(e^{x^{4}}))(e^{x^{2}})-2x(2+\sin(e^{x^{4}}))e^{x^{2}}}{e^{2x^{2}}}$ and we have Proposition 2 for this function. The argument essentially same as that in the proof of Main Theorem 5 completes the proof.
@@
Problem 4 is on singularity theory of differentiable maps and so-called stability of these maps. Stability of smooth maps means that singularities of smooth maps are invariant under slight perturbations. This is a kind of classical theory in singularity theory. For this, see [25, 26, 27, 28, 29, 30, 31, 32], papers on celebrated and sophisticated theory of Mather, and for a textbook on singularity theory of differentiable maps, see [6], for example. For the space of smooth maps from a closed manifold into another smooth manifold with no boundary, so-called stable maps exist densely if the pair of the dimensions of the manifolds is nice in the sense of the presented theory of Mather. For example, if the dimensions are sufficiently low, then the pair is nice in this sense. It has been difficult to find nice criteria to verify stability for non-proper maps, where fundamental criteria have been shown for proper maps. For Morse functions on closed manifolds, a smooth real-valued function on a smooth manifold with no boundary is stable if and only if it is a Morse function whose critical values are distinct at distinct critical values (of the function). Note that there are other kinds of stability of smooth maps, where in the case of proper maps, these are equivalent notions. In [2], Dimca has studied stability of non-proper (Morse) functions on $\mathbb{R}$. In [8], Hayano has studied various kinds of stability of non-proper (Morse) functions on general smooth manifolds with no boundary. For example, in [8], a sufficient condition for a Morse function to be stable is given and stability for some explicit Morse functions are also studied. For recent studies on stability, see also [9, 10]. According to these studies by Ichiki, in non-proper cases, stable maps do not exist densely.
@@
For this, we present an explicit case as Example 4. We omit exposition of the notion of stable functions, strongly stable functions, and infinitesimally stable ones. For rigorous understanding, refer to [8].
@@
  • At distinct critical points of $c$ the values are always distinct.
  • For each point $p\supset c(S(c))$ there exists a small open neighborhood $N_{p}$ in $\mathbb{R}$ and in $c^{-1}(N_{p})$, there exists an open set $V_{p}\subset c^{-1}(N_{p})$ satisfying that $V_{p}\bigcap S(c)$ is empty, that $c^{-1}(N_{p})-V_{p}$ is compact, and $c{\mid}_{V_{p}}$ gives a product bundle over $N_{p}$.
@@
  • At distinct critical points of $c$ the values are always distinct.
@@
In our case, we can check this easily. In addition, in the case $a_{2}>0$ instead, the function is not strongly stable, where it is stable. In any case, the function is not infinitesimally stable, according to general theory of stability and exposition in [8, Subsection 2.2]. More precisely, The restriction of our Morse function ${\pi}_{X_{m,c_{1},c_{2}}}$ to its critical set is non-proper around $0\in\mathbb{R}$ in the target space $\mathbb{R}$.

### 23 Mar 2026 v3

@@ Abstract.
Here, the Reeb space of a smooth function on a manifold with no boundary is fundamental and important. This is the naturally topologized quotient space of the manifold, consisting of all connected components (contours) of the function and is a graph under a certain nice situation. Studies related to the present study were started due to interest of the author in theory of Reeb spaces of so-called non-proper functions. For proper functions, in 2020s, related studies have developed mainly due to Gelbukh and Saeki.
@@
  • $K(C_{p})-C_{p}$ does not contain any point of the form $(c_{i}(x),x)$ ($x\in S(c_{i})$).
@@
A smooth function $c_{p,0}:\mathbb{R}\rightarrow\mathbb{R}$ defined by $c_{p,0}(x):=\frac{2+\sin(e^{x^{2}})}{p(x)}$ enjoys the following properties.
@@
    ${c_{p,0}}^{\prime}(x)=\frac{2xe^{x^{2}}\times\cos(e^{x^{2}})\times{p(x)}-(2+\sin(e^{x^{2}}))\times p^{\prime}(x)}{{p(x)}^{2}}$.
This is a kind of fundamental exercises from elementary calculus. By considering $x$ satisfying $\cos(e^{x^{2}})=\pm 1$ and asymptotic behaviors of $e^{x^{2}}$ and real polynomial functions (or so-called real algebraic functions), we have the following.
@@
Let $c_{p,0}$ be a function of Proposition 1. There exist a sequence $\{r_{{\rm P},j}\}_{j=1}^{\infty}$ of positive numbers diverging to $\infty$ such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm P},j})\}_{j=1}^{\infty}$ diverges to $\infty$ and one such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm P},j})\}_{j=1}^{\infty}$ diverges to -$\infty$. There exist a sequence $\{r_{{\rm N},j}\}_{j=1}^{\infty}$ of negative numbers diverging to -$\infty$ such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm N},j})\}_{j=1}^{\infty}$ diverges to $\infty$ and one such that the sequence $\{{c_{p,0}}^{\prime}(r_{{\rm N},j})\}_{j=1}^{\infty}$ diverges to -$\infty$.
@@
Proposition 3.
A smooth function $c_{p,0,0}:\mathbb{R}\rightarrow\mathbb{R}$ defined by $c_{p,0,0}(x):=\frac{2+\sin(e^{x})}{p(x)}$ enjoys the following properties.
  1. (1)
    The values of the function $c_{p,0,0}$ is always positive. The value $c_{p,0,0}(x)$ converges to 0 as $x$ diverges to $\pm\infty$.
  2. (2)
    The 1st derivative ${c_{p,0,0}}^{\prime}(x)$ of $c_{p,0,0}(x)$ is calculated as
    ${c_{p,0,0}}^{\prime}(x)=\frac{e^{x}\times\cos(e^{x})\times{p(x)}-(2+\sin(e^{x}))\times p^{\prime}(x)}{{p(x)}^{2}}$.
This is also a kind of fundamental exercises from elementary calculus. By considering $x$ satisfying $\cos(e^{x})=\pm 1$ in $x>0$ and asymptotic behaviors of $e^{x}$ and real polynomial functions (or so-called real algebraic functions), we have the following.
Proposition 4.
Let $c_{p,0,0}$ be a function of Proposition 3. There exist a sequence $\{r_{{\rm P},j}\}_{j=1}^{\infty}$ of positive numbers diverging to $\infty$ such that the sequence $\{{c_{p,0,0}}^{\prime}(r_{{\rm P},j})\}_{j=1}^{\infty}$ diverges to $\infty$ and one such that the sequence $\{{c_{p,0,0}}^{\prime}(r_{{\rm P},j})\}_{j=1}^{\infty}$ diverges to -$\infty$. The value ${c_{p,0,0}}^{\prime}(x)$ converges to 0 as $x$ diverges to -$\infty$.
@@
We can also have a case with the properties (1, 3, 4, 6).
  1. (6)
    The sets $S(c_{1})\subset\mathbb{R}$ and $S(c_{2})\subset\mathbb{R}$ are bounded below and unbounded above.
@@
We can define $c_{1}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})$ and $c_{2}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+c_{p,0}(x_{1,2})$, with $c_{p,0}$ from Propositions 1 and 2, and have a case for the first statement. More precisely, we choose $p(x)=x^{2}+1$ for example.
We can define $c_{1}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+\frac{1}{2}\times c_{p,0}(x_{1,2})$ and $c_{2}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+c_{p,0}(x_{1,2})$, with $c_{p,0}$ from Propositions 1 and 2, and have a case for the second statement.
We define $c_{1}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+\frac{1}{2}\times c_{p,0,0}(x_{1,2})$ and $c_{2}(x_{1,2}):=c_{{\rm H},p_{1},p_{2}}(x_{1,2})+c_{p,0,0}(x_{1,2})$, with $c_{p,0,0}$ from Propositions 3 and 4. Remember the asymptotic behavior of $c_{p,0,0}$ and the 1st derivative $c_{p,0,0}$$^{\prime}$ and we can see that the addition of $\frac{1}{2}\times c_{p,0,0}(x_{1,2})$ and that of $c_{p,0,0}(x_{1,2})$ to $c_{{\rm H},p_{1},p_{2}}(x_{1,2})$ do not change the signs of the vaues of the 1st derivatives at the infinity -$\infty$. We can see that this gives a case for the third statement.
@@
We can also have a case with the properties (1, 3, 4, 6).
  1. (6)
    The sets $S(c_{1})\subset\mathbb{R}$ and $S(c_{2})\subset\mathbb{R}$ are bounded below and unbounded above.
@@
We remember Propositions 1 and 2. We define a function of a similar type we can use in the present situation. In short, we define $c_{e,1}(x):=\frac{2+\sin(e^{x^{4}})}{e^{x^{2}}}$ instead. The values of this function are always positive and the value converges to 0 at each infinity. Its 1st derivative is calculated as ${c_{e,1}}^{\prime}(x)=\frac{(4x^{3}e^{x^{4}}\cos(e^{x^{4}}))(e^{x^{2}})-2x(2+\sin(e^{x^{4}}))e^{x^{2}}}{e^{2x^{2}}}$ and we have Proposition 2 for this function.
We remember Propositions 3 and 4. We define a function of a similar type we can use in the present situation. In short, we define $c_{e,2}(x):=\frac{2+\sin(e^{x^{3}})}{e^{x^{2}}}$ instead. The values of this function are always positive and the value converges to 0 at each infinity. Its 1st derivative is calculated as ${c_{e,2}}^{\prime}(x)=\frac{(3x^{2}e^{x^{3}}\cos(e^{x^{3}}))(e^{x^{2}})-2x(2+\sin(e^{x^{3}}))e^{x^{2}}}{e^{2x^{2}}}$ and we have Proposition 4 for this function. Remember the asymptotic behavior of the 1st derivative ${c_{e,2}}^{\prime}(x)$ again.
The argument essentially same as that in the proof of Main Theorem 5 completes the proof.
@@
Problem 4 is on singularity theory of differentiable maps and so-called stability of these maps. Roughly, stability of smooth maps means that singularities of smooth maps are invariant under slight perturbations. This is a kind of classical theory in singularity theory. For this, see [25, 26, 27, 28, 29, 30, 31, 32], papers on celebrated and sophisticated theory of Mather, and for a textbook on singularity theory of differentiable maps, see [6], for example. For the space of smooth maps from a closed manifold into another smooth manifold with no boundary, so-called stable maps exist densely if the pair of the dimensions of the manifolds is nice in the sense of the presented theory of Mather. For example, if the dimensions are sufficiently low, then the pair is nice in this sense. It has been difficult to find nice criteria to verify stability for non-proper maps, where fundamental criteria have been shown for proper maps. For Morse functions on closed manifolds, a smooth real-valued function on a smooth manifold with no boundary is stable if and only if it is a Morse function whose critical values are distinct at distinct critical values (of the function). Note that there are other kinds of stability of smooth maps, where in the case of proper maps, these are equivalent notions. In [2], Dimca has studied stability of non-proper (Morse) functions on $\mathbb{R}$. In [8], Hayano has studied various kinds of stability of non-proper (Morse) functions on general smooth manifolds with no boundary. For example, in [8], a sufficient condition for a Morse function to be stable is given and stability for some explicit Morse functions are also studied. For recent studies on stability, see also [9, 10]. According to these studies by Ichiki, in non-proper cases, stable maps do not exist densely.
@@
For this, we present an explicit case as Example 4. We omit rigirous exposition of the notion of stable functions, strongly stable functions, and infinitesimally stable ones. For rigorous understanding, refer to the original paper [8], with the original papers of Mather, and the textbook [6]. We only present partial or complete characterizations, or necessary, sufficient, or necessary and sufficient conditions for Morse functions to be of these classes.
@@
  • At distinct critical points of $c$ the values are always distinct. Equivalently, the restriction $c{\mid}_{S(c)}$ is injective.
  • For each point $p\supset c(S(c))$, there exist a small open neighborhood $N_{p}$ in $\mathbb{R}$ and an open set $V_{p}$ of $X$ and the following hold.
    • $X-V_{p}$ is compact.
    • $c^{-1}(p)\bigcap V_{p}\bigcap S(c)$ is empty.
    • There exists a diffeomorphism ${\Phi}_{p}:(c^{-1}(p)\bigcap V_{p})\times N_{p}\rightarrow c^{-1}(N_{p})\bigcap V_{p}$ and the function $c{\mid}_{c^{-1}(N_{p})\bigcap V_{p}}\circ{\Phi}_{p}$ gives a smooth trivial bundle over $N_{p}$.
@@
  • The restriction $c{\mid}_{S(c)}$ is injective.
@@
In our case, we can check this easily. In addition, in the case $a_{2}>0$ instead, the function is not strongly stable, where it is stable. In any case, the function is not infinitesimally stable, according to general theory of stability and exposition in [8, Subsection 2.2]. More precisely, the restriction ${\pi}_{m+1,1}{\mid}_{S({\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}})}$ of our Morse function ${\pi}_{m+1,1}{\mid}_{X_{m,c_{1},c_{2}}}$ to its critical set is non-proper around $0\in\mathbb{R}$ in the target space $\mathbb{R}$: it is known as a necessary and sufficient condition for a Morse function $c$ such that $c{\mid}_{S(c)}$ is injective to be infinitesimally stable that $c{\mid}_{S(c)}$ is proper.